# Relative_Constructibility_via_Restricted_ASMs.bib

### v1

@@ Abstract
We restrict the computational power of Gurevich’s abstract state machines to obtain two classes of machine models for generalised computation of sets, namely restricted abstract state machines (RASMs) and restricted abstract state machines with parameters (RASMPs). We derive from each class, a relative computability relation on sets, which is analogous to the Turing reducibility relation on reals. We then prove that the relative computability relation derived from RASMPs is equivalent to relative constructibility.
@@ 1 Introduction
1.1 Generalising Computation
Conventionally, across nearly all fields of study, a computation refers to a terminating run of some computer. In the abstract, modern equivalents of the term computer range from algorithm to program. It should thus be non-contentious to say that, any difficulty in formally defining a computation must come from formally defining a program.
So, what exactly is a program? To a software engineer, it might be a compilable piece of code written in any known programming language, e.g. Java. To a mathematician, it could be any Turing machine. Here, the class of compilable Java codes and the class of Turing machines are two classes defined by different models of computation, at arguably different levels of abstraction. In principle, a model of computation ought to
  • include the definition of a run, as well as
  • specify precise conditions for a run to be terminating.
Following the line of thinking inspired by our two examples, it seems convenient — even natural — to define a program as any object admitted by a fixed model of computation.
Models of computation studied by computer scientists and mathematicians are aplenty. They include common programming languages, along with the semantics (see e.g.[scott] and [meyer]) — in various paradigms — associated with each language. They also include stalwarts from almost a century ago, three of which are
@@ 1.1 Generalising Computation
One obvious commonality of said models of computation is, they admit only finitary programs and runs. It follows that these models forbid infinitary computations, which is understandable and justified, considering infinitary computations are intuitively unrealisable in the real world.
But what if someone takes mathematical and philosophical interest in defining and studying infinitary computations, real-world considerations be damned? Is there inherent technical difficulty in doing so? The answer is, “No, not really,” as evidenced by a multitude of well-received attempts (as much as a subject this niched can be well-received) at generalising computation to the infinite. Such generalisations are appropriately termed models of generalised computation.
Since many decades ago, mathematical logicians have tried their hand at adapting models among (CT1) to (CT3) to allow computations with, and on, arbitrary sets.
From the perspective of recursion theory, where the structure of Turing degrees under Turing reducibility and the jump operator is the main object study, leading candidates include $\alpha$-recursion and $E$-recursion. Both paradigms depend on models of computation that generalises (CT1) with schemata on arbitrary sets (see e.g. [takeuti] and [normann]). These schemata give rise to analogues of important results in classical recursion theory, on sets that can be much larger than reals.
From the perspective of “computer as a machine”, we have Koepke’s ordinal Turing machines (see [koepke1]). Ordinal Turing machines are a generalisation of (CT3): they work like standard Turing machines, except their tapes are now $ORD$-length and they can each refer to finitely many ordinals as parameters.
In any case, the models of generalised computation introduced hitherto rely very much on the Church-Turing thesis for justification of properness. Such justification usually goes along the lines of,
if $X$ nicely generalises a model among (CT1) to (CT3) and the Church-Turing thesis says that the object of generalisation is a model of computation, then $X$ should also be a model of (generalised) computation.
@@
This kind of arguments often lead to unintuitive definitions insofar as computation is concerned, sometimes to the extent of an apology being issued (e.g. [sackserec]). Indeed, modern intuition of computation has evolved to become rather high-level, and the models which motivated the Church-Turing thesis have been relegated to play the roles of convenient mathematical foundations, instead of anything resembling practical mental models of computation.
@@
Perhaps the most celebrated attempt at amalgamating and formalising useful mental models of computations and algorithms, comes in the form of a series of papers (starting with [gurevich]) by Gurevich on abstract state machines. Abstract state machines far more closely resemble “high-level” meta-theoretic representations of Plotkin-style semantics (found e.g. in [plotkin]), than they do anything among (CT1) to (CT3). That these machines have been adapted for use in fields ranging from software development to systems engineering is testament to the intuitiveness of their design. Having said that, how well they are equipped to handle philosophical demands of computing with infinite sets is yet to be vigorously tested, for their applications have so far been finitary in nature.
In this paper, we investigate the promise of abstract state machines in infinitary computations. We make well-meaning restrictions to their definition along the way, so that they are better endowed with our intuition of computation born out of classical computing.
1.2 Generalising Computability
@@
    Universality: the relation should be defined on all sets of ordinals.
@@
    Coherence: if the relative computability of two sets is witnessed under known computational constraints, then the same should be witnessed after lifting these constraints.
The astute reader might ask, “should universality in (D1) not require the inclusion of literally all sets in the domain of the relation?” The truth is, (D1) is phrased this way for convenience of analysis. Note that there is little substantive compromise in this apparently weaker presentation, because every set can be simply encoded into a binary string, according to Proposition 2.2. We shall assume that the designation of (D2) and (D3) as desiderata is intuitive, if not self-evident. Unsurprisingly, each of the well-known models of generalised computation introduced in the previous subsection induces relative computability relation on sets. Let us now briefly check (D1) to (D3) against these relations.
@@
$A\leq_{A}B\text{ iff there is }\alpha\text{ for which }A\leq_{\alpha}B\text{.}$
However, if we interpret $\alpha$ to be a measure of the amount of computational resources available, as is natural, Theodore Slaman’s answer in [slaman] tells us $\leq_{A}$ is doomed to be incoherent. It seems, then, that there is no easy fix to the non-universality of the $\leq_{\alpha}$’s.
On the other hand, the relative computability relation of $E$-recursion, written $\leq_{E}$, is both universal and transitive out of the box. Further, it is common in $E$-recursion to study “runs” restricted to arbitrary $E$-recursively closed initial segment of $L$. Restrictions of this type are computational constraints imposed on the evaluation of $\leq_{E}$. Unfortunately, with respect to these constraints, $\leq_{E}$ is not known to behave coherently (see e.g. “The Divergence-Admissibility Split” on p.11 of [sackserec]).
Taking a leaf from how classical Turing machines were modified to become classical oracle machines, we can straightforwardly augment ordinal Turing machines with set oracles. This allows one to define a relative computability relation $\preceq_{A}$ on arbitrary sets in the same way real oracles are used to define the Turing reducibility relation on the reals. In fact, with reference to $\preceq_{\alpha}$ in Definition 19 of [koepke2],
$A\preceq_{A}B\text{ iff there is }\alpha\text{ for which }A\preceq_{\alpha}B\text{.}$
Such a definition is easily seen to be universal and coherent. It is stated in Theorem 20 of [koepke2] that restricting this relative computability relation to subsets of admissible ordinals $\alpha$ may result in a lack of transitivity. We can however verify that $\preceq_{\alpha}$ is transitive whenever $\alpha$ is regular, and in turn, use it to conclude the transitivity of $\preceq_{A}$.
@@
Appears $\dots$ $\alpha$-recursion $E$-recursion $\alpha$-computability
Table 2: Two universal relative computability relations derived from separate classes of non-universal relations.
@@
Derived $\dots$ the $\leq_{\alpha}$’s the $\preceq_{\alpha}$’s
Unlike the previous models of generalised computation, abstract state machines were not designed with any form of relative computability in mind. In fact, in the name of generality, the definition of an abstract state machine does not even specify the existence of canonical inputs or outputs, which are typically integral to associating a relative computability relation with a computational model. Further, preliminary investigations into the limits of their computational power quickly uncover something philosophically unsatisfying, if not unsettling. We seek to correct this with our restrictions to the definition of abstract state machines.
The way restrictions are added to the notion of abstract state machines results in a new model of generalised computation called restricted abstract state machines with parameters (abbreviated as RASMPs). The main draw of RASMPs is that they possess more concrete guardrails than abstract state machines without compromising too much of the latter’s intuitiveness. Coupled with an input/output paradigm, RASMPs naturally induce a relative computability relation, $\leq^{P}$, on all sets of ordinals. It turns out that $\leq^{P}$ coincides with both $\preceq_{A}$ as well as the relative constructibility relation. Unlike $\preceq_{A}$ though, $\leq^{P}$ tends very much to remain transitive under computational constraints. This makes for interesting comparisons between $\leq^{P}$ and say, the relations in Table 1, when size limitations are imposed.
2 Restricting Abstract State Machines
In this section, we progressively chip away at the very useful, but very inclusive, definition of a sequential abstract state machine, to arrive at a more restrictive machine model which also better captures our intuition of generalised computation.
2.1 Abstract State Machines
Gurevich introduced abstract state machines in [gurevich] as general models of what we may conceptualise as an algorithm. He followed up with a number of collaborations and papers (e.g. [gurevichblass]) on various subclasses of such machines. If we follow the convention of using “algorithm” and “computer program” interchangeably, then abstract state machines are natural candidates for capturing the general notion of computability. However, the claim that such machines are embodiments of algorithms seems to have undergone insufficient scrutiny. We shall start this section by refuting this claim through a set-theoretic argument.
Briefly, a sequential abstract state machine — or sequential ASM (henceforth just ASM) — $\mathfrak{M}$ comprises a class of states $S(\mathfrak{M})$, a class of initial states $I(\mathfrak{M})$, and a class transition function $\tau_{\mathfrak{M}}:S(\mathfrak{M})\longrightarrow S(\mathfrak{M})$ such that
@@ 2.1 Abstract State Machines
    $I(\mathfrak{M})\subset S(\mathfrak{M})$, and
@@
    for all $s\in S(\mathfrak{M})$, $\tau_{\mathfrak{M}}(s)$ has the same base set as $s$, and
@@
    if $s_{1},s_{2}\in S(\mathfrak{M})$ and $s_{1}$ and $s_{2}$ are isomorphic, then $\tau_{\mathfrak{M}}(s_{1})$ and $\tau_{\mathfrak{M}}(s_{2})$ are isomorphic.
@@
For brevity’s sake, some details are withheld here. First, Gurevich originally require states to have only function and constant symbols in their signatures. However, this restriction is purely cosmetic because every relation in a structure can be represented as a boolean function. Second, $\tau$ is supposed to satisfy the bounded exploration postulate, which vaguely means that $\tau$ must perform the same updates to all states with the same interpretations of ground terms in a particular finite set. We shall keep this requirement (or at least its spirit) in mind for this subsection, even if we need not formally verify it. Finally, our definition above can be generalised so that $\tau_{\mathfrak{M}}$ need only be a binary relation – resulting in ASMs which might not be sequential – but that is unnecessary for the point we seek to put across.
Intuitively, a state of an ASM is a configuration of some generalised Turing machine. Each initial state corresponds to an initial configuration, from which one can read off the input (or one may take the initial state itself to be the input). The last state of a terminating run, sometimes called a final state, corresponds in spirit to an end configuration of a Turing machine run, from which one can read off the output. For example, if we have a specific symbol $\ulcorner X\urcorner$ in the signature to represent an analogue of the Turing machine tape, then a terminating run $\vec{r}=(s_{1},\dots,s_{n})$ should give an algorithmic transformation of the input representation $X^{s_{1}}$ into the output representation $X^{s_{n}}$.
We argue against ASMs as a good notion of generalised algorithms. Consider the example below.
@@
Let $\kappa$ be an infinite cardinal. Expand the signature of $(\kappa;\in)$ to include a unary relation symbol $\ulcorner A\urcorner$ and a constant symbol $\ulcorner c\urcorner$. Now if $A_{1}$ and $A_{2}$ are subsets of $\kappa$, then we can define a sequential ASM $\mathfrak{M}(\kappa,A_{1},A_{2})$ as follows:
$s_{1}:=(\kappa;\in,A_{1},0)$
$s_{2}:=(\kappa;\in,A_{2},1)$
@@
$\tau_{\mathfrak{M}(\kappa,A_{1},A_{2})}:=\{(s_{1},s_{2}),(s_{2},s_{2})\}\text{.}$
Observe that $A_{1}$ and $A_{2}$ both interpret $\ulcorner A\urcorner$ on the base set $\kappa$, and that $s_{1}$ is not isomorphic to $s_{2}$ because
  • $s_{1}$ interprets $\ulcorner c\urcorner$ as 00 and $s_{2}$ interprets $\ulcorner c\urcorner$ as $1$, and
  • there is no automorphism of $(\kappa;\in)$ other than the identity on $\kappa$.
This means $S(\mathfrak{M}(\kappa,A_{1},A_{2}))$, $I(\mathfrak{M}(\kappa,A_{1},A_{2}))$ and $\tau_{\mathfrak{M}(\kappa,A_{1},A_{2})}$ satisfy (A1) to (A4) in place of $S(\mathfrak{M})$, $I(\mathfrak{M})$ and $\tau_{\mathfrak{M}}$ respectively. Furthermore, $c$ is specifically added to the signature so that the finite set $\{c\}$ witnesses $\mathfrak{M}(\kappa,A_{1},A_{2})$ satisfy the bounded exploration postulate.
Fix subsets $A_{1}$ and $A_{2}$ of $\kappa$. The only terminating run of $\mathfrak{M}(\kappa,A_{1},A_{2})$ is $(s_{1},s_{2})$. Morally, this is saying that we can algorithmically get from $A_{1}$ to $A_{2}$. If we set $A_{1}$ to be a simply definable set like the empty set, then every subset of $\kappa$ is algorithmically derivable from $A_{1}$. This seems incredible given the information content of arbitrary sets of ordinals, as exemplified by the next proposition.
@@
Proposition 2.2 allows us to code any set as a set of ordinals such that the decoding can be done in an absolute manner. If we believe that such a decoding scheme (particularly, Gödel’s pairing function) is algorithmic, then the definition of ASMs entails that we can algorithmically derive every set in $V$ from just $\emptyset$. Since most set theorists are of the opinion that sets in $V$ can be extremely complicated and far from constructible in a very strong sense, this conclusion is wholly unsatisfactory, if not patently false. In other words, ASMs are too encompassing a paradigm for algorithms, and thus should not be used to draw the line for generalised computability.
@@
$s_{1}:=(H(\kappa);\in,A_{1},0)$
$s_{2}:=(H(\kappa);\in,A_{2},1)\text{.}$
@@
For much of the rest of this section, we will make educated modifications to the definition of an ASM, in an attempt to formulate a philosophically justified notion of algorithms.

### v2

# Relative Constructibility via Generalised Sequential Algorithms

@@ Abstract
We modify Gurevich’s definition of sequential algorithms, so that it becomes amenable to computation with arbitrarily large sets on a sufficiently intuitive level. As a result, two classes of abstract algorithms are obtained, namely generalised sequential algorithms (GSeqAs) and generalised sequential algorithms with parameters (GSeqAPs). We derive from each class a relative computability relation on sets which is analogous to the Turing reducibility relation on reals. We then prove that the relative computability relation derived from GSeqAPs is equivalent to the relative constructibility relation in set theory.
@@ 1 Introduction
Models of computation studied by computer scientists and mathematicians are aplenty. They include common programming languages, along with the semantics (see e.g.[2] and [5]) — in various paradigms — associated with each language. They also include stalwarts from almost a century ago, three of which are
@@
That all these models of computation possess the same computational power is great evidence for the Church-Turing thesis. However, modern intuition of computation has evolved to become rather high-level, and the formal models (CT1) to (CT3) which motivated the Church-Turing thesis have been relegated to play the roles of convenient mathematical foundations, instead of anything resembling practical mental models of computation.
Perhaps the most celebrated attempt at amalgamating and formalising useful mental models of computations was born out of this realisation. In a series of papers (starting with [7]), Gurevich gave (closely related) semantic axiomatisations of abstract algorithms, and showed how these algorithms are captured by sufficiently machine-like models called abstract state machines. Abstract state machines were defined by Gurevich himself way back in the 1980s, although more organised literature thereof only became available in the 1990s. A full definition of these machines can be found in e.g. [6].
@@
Going by the semantic abstraction of Gurevich, algorithms far more closely resemble “high-level” meta-theoretic representations of Plotkin-style semantics (found e.g. in [4]), than they do anything among (CT1) to (CT3). That their syntactic counterparts — the programming languages associated with abstract state machines — have been adapted for use in fields ranging from software development to systems engineering via (what is now known as) the ASM method is testament to the intuitiveness of Gurevich’s design. Historical details regarding the ASM method can be found in [9] and Chapter 9 of [11].
One obvious commonality of all models of computation hitherto mentioned is that they admit only finitary computation, which is understandable and justified, considering infinitary computations are unrealisable in the real world. But are there inherent technical difficulties in formalising infinitary computation? The answer is “No, not really,” as evidenced by a multitude of well-received attempts (as much as a subject this niched can be well-received) at generalising computation to the infinite. Such generalisations are appropriately termed models of generalised computation.
Since many decades ago, mathematical logicians have tried their hand at adapting models among (CT1) to (CT3) to allow computations with and on arbitrary sets.
From the perspective of recursion theory, where the structure of Turing degrees under Turing reducibility and the jump operator is the main object study, leading candidates include $\alpha$-recursion and $E$-recursion. Both paradigms depend on models of computation that generalises (CT1) with schemata on arbitrary sets (see e.g. [1] and [3]). These schemata give rise to analogues of important results in classical recursion theory, on sets that can be much larger than reals.
From the perspective of “computer as a machine”, we have Koepke’s ordinal Turing machines (see [13]). Ordinal Turing machines are a generalisation of (CT3): they work like standard Turing machines, except their tapes are now $ORD$-length and they can each refer to finitely many ordinals as parameters.
In any case, it appears that the models of generalised computation introduced so far rely very much on the Church-Turing thesis for justification of properness. Such justification usually goes along the lines of,
since $X$ nicely generalises a model among (CT1) to (CT3) and the subject of generalisation is a model of computation by the Church-Turing thesis, $X$ should also be a model of (generalised) computation.
This kind of arguments often lead to unintuitive definitions insofar as computation is concerned, sometimes to the extent of an apology being issued (e.g. [16]). In fact, it can be argued that the said models were formulated with specific properties of relative computability — instead of intuitiveness — in mind, to ensure their degrees of computability are equipped with non-trivial structure.
In this paper, we adapt Gurevich’s sequential algorithms for generalised computation. In doing so, we bypass the appeal to the Church-Turing thesis, and provide a “top-down” counterpart to the typical “bottom-up” approach to modelling generalised computation. We make gradual well-meaning modifications to Gurevich’s definition, so that even as our algorithms become more encompassing, the growth in their computational power does not contradict our intuition of computation born out of classical computing. In Subsection 2.1, we see that plainly omitting the bounded exploration postulate from the initial definition of sequential algorithms is wholly unsatisfactory from a set-theoretic standpoint. The remainder of Section 2 is thus geared towards finding the right properties to endow a generalised sequential algorithm with, properties that can plug the gap left behind by the aforementioned omission. Notable milestones in this search/journey include:
  • a very much local, yet infinitary, version of the bounded exploration postulate formulated in Subsection 2.4,
  • a streamlined method of writing the input and reading the output of a terminating run devised in Subsection 2.5,
  • the definition of transfinite runs in Subsection 2.7, accompanied by an argument about the nigh-universality of our method of taking limits, and
  • the addition of ordinal parameters in Subsection 3.2 for a better behaved relative computability relation.
The result of our elaborate adaptation process is a new model of generalised computation called generalised sequential algorithms with parameters (abbreviated as GSeqAPs). The first draw of GSeqAPs is that their more structured templates facilitate the design of generalised algorithms, without compromising too much of the intuitiveness in Gurevich’s conception. Coupled with an input/output paradigm, GSeqAPs naturally induce a relative computability relation, $\leq^{P}$, on sets of ordinals. It turns out that $\leq^{P}$ possesses a particular suite of desirable traits not found in other relative computability relations associated with generalised computation, especially under computational constraints (Table 3, as well as the following subsection). This is the second draw of GSeqAPs.
Towards a generalised analogue of the Church-Turing thesis, we also show that $\leq^{P}$ coincides with the relative constructibility relation (Theorem 3.41).
1.1 Generalisations of (Relative) Computability
@@
    Universality: the relation should be reflexive on all sets of ordinals.
@@
    Coherence: if the relative computability of two sets is witnessed under reasonable computational constraints, then the same should be witnessed after lifting these constraints.
Unsurprisingly, each of the well-known models of generalised computation introduced in the previous subsection induces relative computability relation on sets. We shall examine whether these relations satisfy properties (D1) to (D3). The next three paragraphs contain an elaboration of our examination, and Tables 1 and 2, a summary of its outcomes.
@@
[…]
However, if we interpret $\alpha$ to be a measure of the amount of computational resources available, as is natural, Theodore Slaman’s answer in [18] tells us $\leq_{A}$ is doomed to be incoherent. It seems, then, that there is no easy fix to the non-universality of the $\leq_{\alpha}$’s.
On the other hand, the relative computability relation of $E$-recursion, written $\leq_{E}$, is both universal and transitive out of the box. Further, it is common in $E$-recursion to study “runs” restricted to arbitrary $E$-recursively closed initial segment of $L$. Restrictions of this type are computational constraints imposed on the evaluation of $\leq_{E}$. Unfortunately, with respect to these constraints, $\leq_{E}$ is not known to behave coherently (see e.g. “The Divergence-Admissibility Split” on p.11 of [16]).
Taking a leaf from how classical Turing machines were modified to become classical oracle machines, we can straightforwardly augment ordinal Turing machines with set oracles. This allows one to define a relative computability relation $\preceq_{A}$ on arbitrary sets in the same way real oracles are used to define the Turing reducibility relation on the reals. In fact, with reference to $\preceq_{\alpha}$ in Definition 19 of [14],
[…]
Such a definition is easily seen to be universal and coherent. It is stated in Theorem 20 of [14] that restricting this relative computability relation to subsets of admissible ordinals $\alpha$ may result in a lack of transitivity. We can however verify that $\preceq_{\alpha}$ is transitive whenever $\alpha$ is regular, and in turn, use it to conclude the transitivity of $\preceq_{A}$.
@@
[…]
Table 2: Comparison of two universal relative computability relations derived from separate classes of non-universal relations.
@@
[…]
Like $\preceq_{A}$, the relative computability relation associated with GSeqAPs, $\leq^{P}$, also fulfils desiderata (D1) to (D3). What makes $\leq^{P}$ stand out is that unlike $\preceq_{A}$, it tends very much to remain transitive under computational constraints. More precisely, for all admissible ordinals $\alpha$, both
  • $\leq^{P}_{\alpha}$, the restriction of $\leq^{P}$ to computations requiring $\alpha$-sized space, and
  • $\leq^{P,s}_{\alpha}$, the restriction of $\leq^{P}$ to computations requiring $\alpha$-sized space and terminating in $\alpha-$many time steps,
are at once, transitive and coherent (Propositions 3.11, 3.12, 3.27 and 3.28). It is interesting to note that our formulation of GSeqAPs, in accordance to Gurevich’s principles of inclusiveness and intuitiveness espoused in [7], inadvertently gives rise to a series of novel, non-trivial and well-behaved relative computability relations.
2 Adapting Sequential Algorithms for the Infinite
In this section, we progressively modify the definition of a sequential algorithm, to arrive at a more concrete notion of algorithms which also better captures our intuition of generalised computation.
2.1 Formal Sequential Algorithms
Gurevich introduced the ASM thesis in [7] as a justification of correctness for his definition of (a most general conceptualisation of) a sequential algorithm. He followed it up with a number of collaborations and papers (e.g. [10]) on various other classes of algorithms. If we follow the convention of using “algorithm” and “computer program” interchangeably, then sequential algorithms are natural candidates for capturing a general notion of computability. However, it seems that Gurevich himself did not intend for sequential algorithms to capture computation on an infinitary scale. We shall start this section with a brief examination on how his definition might be ill-equipped at dealing with generalised computation.
Briefly, a sequential algorithm — henceforth SeqA, to distinguish the formal concept from the colloquial — $\mathfrak{M}$ comprises a class of states $S(\mathfrak{M})$, a class of initial states $I(\mathfrak{M})$, a class transition function $\tau_{\mathfrak{M}}:S(\mathfrak{M})\longrightarrow S(\mathfrak{M})$, and a finite set $G$ of ground terms over $\sigma(\mathfrak{M})$ such that, letting $G^{s}$ denote the function $\dot{x}\mapsto\dot{x}^{s}$ with domain $G$,
@@ 2.1 Formal Sequential Algorithms
    $I(\mathfrak{M})\subset S(\mathfrak{M})$,
@@
    for all $s\in S(\mathfrak{M})$, $\tau_{\mathfrak{M}}(s)$ has the same base set $b_{\mathfrak{M}}(s)$ as $s$
@@
    for all $s\in S(\mathfrak{M})$, $\ulcorner X\urcorner\in\sigma(\mathfrak{M})$, $n<\omega$ and $\vec{a}\in b_{\mathfrak{M}}(s)^{n}$, if $n$ is the arity of $\ulcorner X\urcorner$, then
    $\ulcorner X\urcorner^{s}(\vec{a})\neq\ulcorner X\urcorner^{\tau_{\mathfrak{M}}(s)}(\vec{a})\implies\vec{a}^{\frown}(\ulcorner X\urcorner^{\tau_{\mathfrak{M}}(s)}(\vec{a}))\in ran(G^{s})^{n+1}$
    whenever $\ulcorner X\urcorner$ is a function or constant symbol and
    $\neg(\ulcorner X\urcorner^{s}(\vec{a})\iff\ulcorner X\urcorner^{\tau_{\mathfrak{M}}(s)}(\vec{a}))\implies\vec{a}\in ran(G^{s})^{n}$
    whenever $\ulcorner X\urcorner$ is a relation symbol, and
  5. (A5)
    for all $s,s^{\prime}\in S(\mathfrak{M})$, $\ulcorner X\urcorner\in\sigma(\mathfrak{M})$ and $n<\omega$, if $n$ is the arity of $\ulcorner X\urcorner$ and $G^{s}=G^{s^{\prime}}$, then whenever $\vec{a}\in ran(G^{s})^{n}$,
    $\ulcorner X\urcorner^{s}\text{ agrees with }\ulcorner X\urcorner^{\tau_{\mathfrak{M}}(s)}\text{ at }\vec{a}\iff\$$\ulcorner X\urcorner^{s^{\prime}}\text{ agrees with }\ulcorner X\urcorner^{\tau_{\mathfrak{M}}(s^{\prime})}\text{ at }\vec{a}\text{, and}$
    $\ulcorner X\urcorner^{s}\text{ disagrees with }\ulcorner X\urcorner^{\tau_{\mathfrak{M}}(s)}\text{ at }\vec{a}\implies\$$\ulcorner X\urcorner^{\tau_{\mathfrak{M}}(s)}\text{ agrees with }\ulcorner X\urcorner^{\tau_{\mathfrak{M}}(s^{\prime})}\text{ at }\vec{a}\text{.}$
@@
For brevity’s sake, some details are withheld here. First, Gurevich originally require states to have only function and constant symbols in their signatures. However, this restriction is purely cosmetic because every relation in a structure can be represented as a boolean function. Second, $S(\mathfrak{M})$ is supposed to be closed under all possible isomorphisms, thus becoming a proper class. This closure property allows one to derive (A4) and (A5) from Gurevich’s bounded exploration postulate (see [7], Lemma 6.2 for a derivation). Even though our presentation might seem logically weaker than Gurevich’s original, the two are actually provably equivalent, modulo closure under isomorphisms. In fact, going through (A4) and (A5) is standard for proofs equating the computational power of different paradigms through sequential algorithms (in e.g. [15] or even [7] itself), so our sidestepping of set-theoretic concerns regarding the sizes of state spaces is accompanied by some conveniences. Finally, our definition above can be generalised so that $\tau_{\mathfrak{M}}$ need only be a binary relation – resulting in algorithms which might not be sequential – but that is unnecessary if we only want to capture enough of generalised computation for a sufficiently general relative computability relation.
Intuitively, a state of an algorithm is a configuration of some generalised Turing machine. Each initial state corresponds to an initial configuration, from which one can read off the input (or one may take the initial state itself to be the input). The last state of a terminating run, sometimes called a final state, corresponds in spirit to an end configuration of a Turing machine run, from which one can read off the output. For example, if we have a specific symbol $\ulcorner X\urcorner$ in the signature to represent an analogue of the Turing machine tape, then a terminating run $(s_{1},\dots,s_{n})$ should give an algorithmic transformation of the input representation $X^{s_{1}}$ into the output representation $X^{s_{n}}$.
Now, requirements (A4) and (A5) set an a priori finite bound to the number of bits that an algorithm can alter in each step of its computational run. Since a terminating run is presumed to be finite, there is no way for an algorithm to make an infinite amount of change before terminating. It is therefore very tempting to just free ourselves from the shackles of (A4) and (A5), especially since we aim to have as general a definition of an generalised algorithm as possible. Unfortunately, this turns out to be too inclusive. Consider the example below, modelled after a Turing machine with a $\kappa$-length tape.
@@
Let $\kappa$ be an infinite cardinal. Expand the signature of $(\kappa;\in)$ to include a unary relation symbol $\ulcorner A\urcorner$. Now if $A_{1}$ and $A_{2}$ are subsets of $\kappa$, then we can define $\mathfrak{M}(\kappa,A_{1},A_{2})$ as follows:
$s_{1}:=(\kappa;\in,A_{1})$
$s_{2}:=(\kappa;\in,A_{2})$
@@
$\tau_{\mathfrak{M}(\kappa,A_{1},A_{2})}:=\{(s_{1},s_{2}),(s_{2},s_{2})\}\text{,}$
so that $A_{1}$ and $A_{2}$ both interpret $\ulcorner A\urcorner$ on the base set $\kappa$ shared between the states $s_{1}$ and $s_{2}$.
Fix subsets $A_{1}$ and $A_{2}$ of $\kappa$. Note that $\mathfrak{M}(\kappa,A_{1},A_{2})$ satisfies requirements (A1) to (A3). Moreover, the only terminating run of $\mathfrak{M}(\kappa,A_{1},A_{2})$ is $(s_{1},s_{2})$. As the interpretation of the symbol $\ulcorner A\urcorner$ is the only difference between $s_{1}$ and $s_{2}$, the moral implication of this run is that we can algorithmically get from $A_{1}$ to $A_{2}$. If we set $A_{1}$ to be a simply definable set like the empty set, then every subset of $\kappa$ is algorithmically derivable from $A_{1}$. This seems incredible given the information content of arbitrary sets of ordinals, as exemplified by the next proposition.
@@
The procedure described in the proof of Proposition 2.2 allows us to encode any set as a set of ordinals, such that the decoding can be done in an absolute manner. If we believe that such a decoding scheme (particularly, Gödel’s pairing function) is algorithmic, then our aforementioned loosening of the definition of SeqAs entails that we can algorithmically derive every set in $V$ from just $\emptyset$. Since most set theorists are of the opinion that sets in $V$ can be extremely complicated and far from constructible in a very strong sense, this conclusion is wholly unsatisfactory, if not patently false. In other words, just eschewing requirements (A4) and (A5) yields too encompassing a paradigm for algorithms, and thus should not be used to draw the line for generalised computability.
@@
$s_{1}:=(H(\kappa);\in,A_{1})$
$s_{2}:=(H(\kappa);\in,A_{2})\text{.}$
@@
From now on, and until further notice, let us relax the definition of SeqAs such that they need only satisfy requirements (A1) to (A3). For much of the rest of this section, we will make educated modifications and impose suitable restrictions to this relaxed definition of a SeqA, in an attempt to formulate a more philosophically justified notion of generalised algorithms, at least from a set-theoretic viewpoint.